# Good Modelling Software Practices

### v2

@@ 1 Introduction
Frequently in socio-environmental sciences, models are used as tools to represent, understand, project and predict the behaviour of these complex systems, and for many more purposes (Epstein 2008; Edmonds et al. 2019). The degree of a model’s formalisation ranges from conceptual to mathematical equations to implementation in software (Romanowska 2015), and–by definition–all of these models are purpose-driven simplifications of the system they represent (Stachowiak 1973). We here concentrate on computational models, i.e. on socio-environmental models implemented in software, and there are many of those out there: Currently the CoMSES Net (2024) lists 1153 agent-based models (ABM); two decades ago, there were 1360 ecological models (Benz, Hoch, and Legović 2001); in 2015, a survey amongst 42 modellers counted 278 aquatic ecosystem models (A. B. G. Janssen et al. 2015).
@@
GMP appear as five phases in the model life cycle (Table 1 of Jakeman et al. 2024): (1) Problem scoping, (2) conceptualisation, (3) formulation and evaluation, (4) application, and (5) perpetuation, and the reiteration of these phases. Good Modelling Software Practices (GMSP) are prominent in phases 3–5 and are thus subsumed under GMP. But where GMP is concerned with the reliability and validity of a model—possibly foremost its —GMSP is concerned with the quality and reliability of a model’s software implementation and beyond: a good quality model software is not restricted to good computer code, but will support iterations through the model life cycle and will enable Good Scientific Practice (DFG 2013; All European Academies 2023).
@@
We start off by motivating each of the good practices and contextualizing them towards the goal of publishing a model software or a scientific result arising from it; we put GMSP at the center of a triangle formed by good practices in modelling, software development, and research (Figure 1). We describe the tools that can be used in a non-exhaustive way covering the entire range of good practices; you may deviate from the tools we selected, or disagree with them, and you may add others or leave some out that we suggested.
@@ 3.1 Version control software
A transparent and reproducible distributed SCM (formerly version control system, VCS) is the basis for good software and has been termed “possibly the biggest advance in software development technology” (Spolsky 2010). The currently dominant SCM software is Git [^2] https://git-scm.com OSS distributed VCS. Note: all Uniform Resource Locators (URL) in this manuscript have been last visited and checked on August 20, 2024, originally invented by Linus Torvalds, the creator of Linux. As source code is text, the SCM tracks changes in lines or parts of lines of text. It can be very well used to manage other kinds of changing texts, such as the text of this manuscript. In fact, this manuscript was started with git init; git add manuscript.md; git commit -m "feat: Created manuscript". Some care should be taken with the latter commit messages, as they should be short and descriptive to humans, and at best also machine-interpretable, e.g., by following the conventional commits [^3] https://www.conventionalcommits.org How to add human and machine readable commit meaning recommendations. Learning Git could be considered a valuable investment in yourself for any creative work.
@@ 3.2 Source code management service
The most prominent online SCM service is GitHub [^6] https://github.com Public GitHub, but many academic institutions also offer on-premise or cross-institutional dedicated SCM services, such as the community GitLab of the German Helmholtz Association [^7] https://codebase.helmholtz.cloud.de Community Gitlab of the German Helmholtz Association, for their students and researchers. A good reason to choose GitHub is the higher number of potential contributors on this platform, estimated at roughly 100 million. Many automated tools for supporting software development only work on GitHub. On the other hand, on-premise GitLabs may be preferred by academic institutions, because the code is then hosted in the research centre or by a dedicated subcontracted partner and may offer better data and IP protection; it may also offer more elaborated services or dedicated computing resources.
@@ 3.3.2 Collaboration
Traditionally in science, mostly the authors of scientific publications are acknowledged and cited. But for scientific software, there are more roles that do not qualify for authorship: foremost those of the research software engineers, but also, e.g., data managers, administrators, friendly reviewers, HPC support staff. The All Contributors [^20] https://allcontributors.org Contributor acknowledgement automation bot integrates with the GitHub SCM service and facilitates proper acknowledgement of those and other non-author groups.
@@ 3.5 Bundling your application
Much of research software, and especially those (like climate models) with a long history, are distributed by giving access to the repository hosted on an online SCM service, or to an archive file that contains the entire source code. Users are then expected to install the (often numerous) requirements, and eventually compile the code and run the model. To improve usability, more state-of-the art software engineering techniques like packaging and containerisation are available (Vanegas Ferro et al. 2022). Not only do they standardise and simplify the installation of the code, but they also improve its findability via machine- and human-readable metadata, support reusability via versioning and ensure proper citeability. We refer to this as bundling your application. This bundling can be integrated in the CD by deploying the bundle in the associated registries of the online SCM services [^24] https://docs.github.com/en/packages, [^25] https://docs.gitlab.com/ee/user/packages/. You can distribute a package of your model software or your model software including a runtime environment
@@ 3.5.2 Container images
Installation can further be simplified for the user by distributing the software as a container image. Like packages, container images are distributed as a single file. They contain an entire operating system layer where the necessary software and its dependencies have been installed already. These images can then be published in container registries to make them reusable. A prominent tool to build containers is Docker [^29] https://www.docker.com/ Docker tool to build containers, a prominent registry the Docker Hub [^30] https://hub.docker.com/ Free to use container image registry, or the built-in container registries of the online SCM services GitHub and GitLab [^31] https://docs.github.com/en/packages, [^32] https://docs.gitlab.com/ee/user/packages/. Apptainer [^33] https://apptainer.org HPC container builder (formerly Singularity) provides container imaging for for High Performance Computing (HPC).
@@ 3.7.1 Code formatting and linting
Automated formatters help adhere to conventions. Tools like Black [^39] https://black.readthedocs.io Python formatter or isort [^40] https://pycqa.github.io/isort/ Python import sorter adjust (Python) code to meet specific formatting guidelines, eliminating the need for manual formatting and ensuring consistency across the codebase. A general-purpose tool for many file formats and languages is Prettier [^41] https://prettier.io General purpose code prettifier.
@@ 3.8 Archiving
Archiving the source code of the model and post-processing routines is essential Good Scientific Practice (DFG 2013). The simplest solution for open-source projects is using the Software Heritage Project [^46] https://www.softwareheritage.org Software heritage archive crawler. Many SCM services are already crawled by this project, so when you upload your public repository, it will be archived automatically; you can also add your project manually.
@@ 4 Good enough modelling software practice - a use case
For reproducibility, a conda environment-dev.yaml for satisfying developer dependencies, as well as a pyproject.toml for the python dependencies, are provided. The model has been archived on Zenodo, and on CoMSES.net together with its ODD (Lemmen et al. 2023). A software paper has been published in JOSS (Lemmen et al. 2024) with a transparent review process [^56] https://github.com/openjournals/joss-reviews/issues/5731; the model has been catalogued in the ABM specific CoMSES.net as well as the Helmholtz research software directory.
@@ 5 Conclusion
Good Practices have been maturing in the areas of software, modelling, and research. Adopting all of them for socio-environmental modelling establishes Good Modelling Software Practices: Reproducibility and citeability meet automated verification, rights management and collaboration meet purpose-driven simplification and validation. For this, you must have version control and distributed redundancy, use a collaboration platform, have a license and documentation, provide readable code, and archive and publish your model software. You should use CI, have guidelines for collaboration, elaborate on documentation, employ versioning and releases, use automated tools for formatting, linting, and verification and target a community platform. You may use CD and packages (Figure 2).

### v4

@@ 1 Introduction
Frequently in socio-environmental sciences, models are used as tools to represent, understand, project and predict the behaviour of these complex systems, and for many more purposes (Epstein 2008; Edmonds et al. 2019). The degree of a model’s formalisation ranges from conceptual to mathematical equations to implementation in software (Romanowska 2015), and–by definition–all of these models are purpose-driven simplifications of the system they represent (Stachowiak 1973). We here concentrate on computational models, i.e. on socio-environmental models implemented in software, and there are many of those out there: Currently the CoMSES Net (2024) lists 1153 agent-based models (ABM, see list of abbreviations provided); two decades ago, there were 1360 ecological models (Benz, Hoch, and Legović 2001); in 2015, a survey amongst 42 modellers counted 278 aquatic ecosystem models (A. B. G. Janssen et al. 2015).
@@
GMP appear as five phases in the model life cycle (Jakeman et al. 2024, their Figure 1): (1) Problem scoping, (2) conceptualisation, (3) formulation and evaluation, (4) application, and (5) perpetuation, and the reiteration of these phases. Good Modelling Software Practices (GMSP) are prominent in phases 3–5 and are thus subsumed under GMP. But where GMP is concerned with the reliability and validity of a model, GMSP is concerned with the quality and reliability of a model’s software implementation and beyond: a good quality model software is not restricted to good computer code, but will support iterations through the model life cycle and will enable Good Scientific Practice (DFG 2013; All European Academies 2023).
@@
We start off by motivating each of the good practices and contextualizing them towards the goal of publishing a model software or a scientific result arising from it; we put GMSP at the center of a triangle formed by good practices in modelling, software development, and research (Figure 1); we categorize broadly which practices must, which ones should and which ones may be implemented, summarized in Figure 2. We describe the tools that can be used in a non-exhaustive way covering the entire range of good practices; you may and will deviate from the tools we selected, or disagree with them, and you may add others or leave some out that we suggested.
@@ 3.1 Version control software
A transparent and reproducible distributed SCM (formerly version control system, VCS) is the basis for good software and has been termed “possibly the biggest advance in software development technology” (Spolsky 2010). The currently dominant SCM software is Git [^2] https://git-scm.com OSS distributed VCS. Note: all Uniform Resource Locators (URL) in this manuscript have been last visited and checked on September 16, 2024, originally invented by Linus Torvalds, the creator of Linux. As source code is text, the SCM tracks changes in lines or parts of lines of text. It can be very well used to manage other kinds of changing texts, such as the text of this manuscript. In fact, this manuscript was started with git init; git add manuscript.md; git commit -m "feat: Created manuscript". Some care should be taken with the latter commit messages, as they should be short and descriptive to humans, and at best also machine-interpretable, e.g., by following the conventional commits [^3] https://www.conventionalcommits.org How to add human and machine readable commit meaning recommendations. Learning Git could be considered a valuable investment in yourself for any creative work.
@@ 3.2 Source code management service
The most prominent online SCM service is GitHub [^6] https://github.com Public GitHub SCM service, but many academic institutions also offer on-premise or cross-institutional dedicated SCM services, such as the community GitLab of the German Helmholtz Association [^7] https://codebase.helmholtz.cloud Community Gitlab of the German Helmholtz Association, for their students and researchers. A good reason to choose GitHub is the higher number of potential contributors on this platform, estimated at roughly 100 million. Many automated tools for supporting software development only work on GitHub. On the other hand, on-premise GitLabs may be preferred by academic institutions, because the code is then hosted in the research centre or by a dedicated subcontracted partner and may offer better data and IP protection; it may also offer more elaborated services or dedicated computing resources.
@@ 3.3.2 Collaboration
Traditionally in science, mostly the authors of scientific publications are acknowledged and cited. But for scientific software, there are more roles that do not qualify for authorship: foremost those of the research software engineers, but also, e.g., data managers, administrators, friendly reviewers, high-performance computing (HPC) support staff. The All Contributors [^20] https://allcontributors.org Contributor acknowledgement automation bot integrates with the GitHub SCM service and facilitates proper acknowledgement of those and other non-author groups.
@@ 3.5 Bundling your application
Much of research software, and especially those (like climate models) with a long history, are distributed by giving access to the repository hosted on an online SCM service, or to an archive file that contains the entire source code. Users are then expected to install the (often numerous) requirements, and eventually compile the code and run the model. To improve usability, more state-of-the art software engineering techniques like packaging and containerisation are available (Vanegas Ferro et al. 2022). Not only do they standardise and simplify the installation of the code, but they also improve its findability via machine- and human-readable metadata, support reusability via versioning and ensure proper citeability. We refer to this as bundling your application. This bundling can be integrated in the CD by deploying the bundle in the associated registries of the online SCM services [^24] https://docs.github.com/en/packages, [^25] https://docs.gitlab.com/ee/user/packages. You can distribute a package of your model software or your model software including a runtime environment
@@ 3.5.2 Container images
Installation can further be simplified for the user by distributing the software as a container image. Like packages, container images are distributed as a single file. They contain an entire operating system layer where the necessary software and its dependencies have been installed already. These images can then be published in container registries to make them reusable. A prominent tool to build containers is Docker [^29] https://www.docker.com Docker tool to build containers, a prominent registry the Docker Hub [^30] https://hub.docker.com Free to use container image registry, or the built-in container registries of the online SCM services GitHub and GitLab [^31] https://docs.github.com/en/packages, [^32] https://docs.gitlab.com/ee/user/packages. Apptainer [^33] https://apptainer.org HPC container builder (formerly Singularity) provides container imaging for for High Performance Computing (HPC).
@@ 3.7.1 Code formatting and linting
Automated formatters help adhere to conventions. Tools like Black [^39] https://black.readthedocs.io Python formatter or isort [^40] https://pycqa.github.io/isort Python import sorter adjust (Python) code to meet specific formatting guidelines, eliminating the need for manual formatting and ensuring consistency across the codebase. A general-purpose tool for many file formats and languages is Prettier [^41] https://prettier.io General purpose code prettifier.
@@ 3.8 Archiving
Archiving the source code of the model and post-processing routines is essential Good Scientific Practice (DFG 2013). The simplest solution for open-source projects is using the Software Heritage Project [^46] https://www.softwareheritage.org Software archive crawler. Many SCM services are already crawled by this project, so when you upload your public repository, it will be archived automatically; you can also add your project manually.
@@ 4 Good enough modelling software practice - a use case
For reproducibility, a conda environment-dev.yaml for satisfying developer dependencies, as well as a pyproject.toml for the python dependencies, are provided. The model has been archived on Zenodo, and on CoMSES.net together with its ODD (Lemmen et al. 2023). A software paper has been published in JOSS (Lemmen et al. 2024) with a transparent review process [^56] https://github.com/openjournals/joss-reviews/issues/5731 Transparent review; the model has been catalogued in the ABM specific CoMSES.net as well as the Helmholtz research software directory.
@@ 5 Conclusion
Good Practices have been maturing in the areas of software, modelling, and research. Adopting all of them for socio-environmental modelling establishes Good Modelling Software Practices: Reproducibility and citeability meet automated verification, rights management and collaboration meet purpose-driven simplification and validation. For this, you must have version control and distributed redundancy, use a collaboration platform, have a license and documentation, provide readable code, and archive and publish your model software. You should use CI, have guidelines for collaboration, elaborate on documentation, employ versioning and releases, use automated tools for formatting, linting, and verification and target a community platform. You may use CD and packages.